OpenStudy (anonymous):

6-(-81)=? my mind went blank help please

hartnn (hartnn):

2 negatives become positive \(6-(-81)=6+81=...?\)

OpenStudy (anonymous):

omg you are awesome

OpenStudy (anonymous):

oh thank you

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

jup